### v1

# On finite generating sets of infinitely generated ideals

Abstract.
This paper presents a novel approach to constructing finite generating sets for infinitely generated ideals. By integrating algebraic and computational techniques, we provide a method to identify finite generators, demonstrated through illustrative examples.
2020 Mathematics Subject Classification:
Primary 13E15; Secondary 13P10.
1. Introduction
The Hilbert basis theorem [Hilbert] asserts that any polynomial ideal is finitely generated. Nonetheless, determining a finite generating set for an ideal derived from infinitely many polynomials poses significant theoretical and computational challenges. For example, take the ideal of the polynomial ring ${\mathbb{Q}}[x,y,z,w]$ generated by infinitely many polynomials
$f_{n}=(2n^{6}+3)xy^{n^{3}+2}+(n^{3}+1)z^{n^{3}+2}+w^{n^{3}+2},n\in{\mathbb{Z}}_{\geq 0}.$
Later, we will demonstrate that this ideal is finitely generated by $f_{0},f_{1},f_{2},f_{3}$ (Example 3.2).
We first consider ideals generated by polynomials $f_{n_{1},\dots,n_{p}}$ where the parameters $n_{1},\dots,n_{p}$ appear in the coefficients and exponents in polynomial forms, and range over the non-negative integers. However, this general form of the problem is not solvable even in the case of a polynomial ring with a single variable: Let $\varphi(n_{1},\dots,n_{p})\in{\mathbb{Z}}[n_{1},\dots,n_{p}]$ be a polynomial, and consider the following ideal of ${\mathbb{Q}}[x]$
$I=\langle x^{\varphi(n_{1},\dots,n_{p})^{2}}\mid n_{1},\dots,n_{p}\in{\mathbb{Z}}_{\geq 0}\rangle\subset{\mathbb{Q}}[x].$
Then $I$ is generated by $x^{M}$ where $M=\min\{\varphi(n_{1},\dots,n_{p})^{2}\mid n_{1},\dots,n_{p}\in{\mathbb{Z}}_{\geq 0}\}$. In particular, $I={\mathbb{Q}}[x]$ if and only if the Diophantine equation $\varphi(n_{1},\dots,n_{p})=0$ has a non-negative integer solution. Therefore, determining a generating set of $I$ depends on the solvability of the Diophantine equation $\varphi(n_{1},\dots,n_{p})=0$. According to the MRDP (Matiyasevich-Robinson-Davis-Putnam) theorem [Matiyasevich], no algorithm exists to decide this in general.
Therefore, constraints must be imposed on the exponents to facilitate the derivation of finite generating sets. Beginning with the simplest case where the exponents are linear, we provide explicit methods to identify finite generating sets.
Main Theorem (Theorem 2.5).
Fix positive integers $p,r$, and $N$. Let $R$ be a commutative ring, and $x_{1},\dots,x_{r}\in R$. Let
$c_{k}(n_{1},\dots,n_{p})\in R[n_{1},\dots,n_{p}],1\leq k\leq N,$
be polynomials in the parameters $n_{1},\dots,n_{p}$, and let $\alpha^{(k)}_{ij},\beta^{(k)}_{i}\in{\mathbb{Z}}_{\geq 0}$ be non-negative integers for $1\leq i\leq r,1\leq j\leq p,1\leq k\leq N$. Let
$f_{n_{1},\dots,n_{p}}=\sum_{k=0}^{N}c_{k}(n_{1},\dots,n_{p})\prod_{i=1}^{r}x_{i}^{\sum_{j=1}^{p}\alpha^{(k)}_{ij}n_{j}+\beta^{(k)}_{i}},$
and consider the ideal $I=\langle f_{n_{1},\dots,n_{p}}\mid{n_{1},\dots,n_{p}}\in{\mathbb{Z}}_{\geq 0}\rangle\subset R$. We set
$\ell_{i}=\sum_{k=1}^{N}(\mathop{\mathrm{deg}}\nolimits_{n_{i}}c_{k}+1)-1.$
Then $I$ is generated by $\{f_{n_{1},\dots,n_{p}}\mid\forall i,0\leq n_{i}\leq\ell_{i}\}$.
In the last section, we will show how to apply the above theorem along with the Gröbner basis theory in the case of higher degree exponents.
2. Main theorem
Let $R$ be a commutative ring (not necessarily Noetherian). For $\bm{x}=(x_{1},\dots,x_{r})\in R^{r}$ and an exponent vector $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{r})\in{\mathbb{Z}}_{\geq 0}^{r}$, we use the multi-index notation $\bm{x^{\alpha}}=x_{1}^{\alpha_{1}}\cdots x_{r}^{\alpha_{r}}$. For example, let $R={\mathbb{Q}}[x,y]$, $\bm{z}=(2,x,x+y)\in R^{3}$, $\bm{\alpha}=(n+2,n,2n+1)$, and $\bm{\beta}=(n,0,0)$, then $\bm{z}^{\bm{\alpha}}+\bm{z}^{\bm{\beta}}=2^{n+2}x^{n}(x+y)^{2n+1}+2^{n}$.
We define the product ordering $\leq$ on ${\mathbb{Z}}_{\geq 0}^{r}$ as
$\bm{\alpha}=(\alpha_{1},\dots,\alpha_{r})\leq\bm{\beta}=(\beta_{1},\dots,\beta_{r})\Leftrightarrow 1\leq\forall i\leq r,\alpha_{i}\leq\beta_{i}.$
For example, $(2,7,4)\leq(3,7,6)$, and $(1,3,1)$ and $(4,1,4)$ are not comparable.
$R\llbracket\bm{t}\rrbracket=R\llbracket t_{1},\dots,t_{r}\rrbracket$ denotes the formal power series ring over $R$. For a infinite collection $\{f_{\bm{n}}\in R\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r}\}$, we consider its generating function $\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r}}f_{\bm{n}}\bm{t}^{\bm{n}}$. Note that $\psi(\bm{t})\in R\llbracket\bm{t}\rrbracket$ is invertible if and only if $\psi(\bm{0})$ is invertible in $R$.
The next lemma is the key to prove the main theorem..
Lemma 2.1.
Fix a vector $\bm{\beta}_{0}=(\beta_{1},\dots,\beta_{r})\in{\mathbb{Z}}_{\geq 0}^{r}$. Let $f_{\bm{n}}\in R$ for $\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r}$, and $g_{\bm{\alpha}}\in R$ for $\bm{\alpha}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{\alpha}\leq\bm{\beta}_{0}$. If there exists an invertible element $\psi(\bm{t})\in R\llbracket\bm{t}\rrbracket$ such that
$\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r}}f_{\bm{n}}\bm{t}^{\bm{n}}=\psi(\bm{t})\sum_{\bm{\alpha}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{\alpha}\leq\bm{\beta}_{0}}g_{\bm{\alpha}}\bm{t}^{\bm{\alpha}},$
then the ideal $\langle f_{\bm{n}}\in R\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r}\rangle\subset R$ is generated by the finite set
$\{f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{n}\leq\bm{\beta}_{0}\}.$
Proof.
By expanding the right-hand side and comparing the coefficients on both sides of the equation, we conclude that $f_{\bm{n}}\in\langle g_{\bm{\alpha}}\mid\bm{\alpha}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{\alpha}\leq\bm{\beta}_{0}\rangle$ for all $\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r}$. Thus $\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r}\rangle\subset\langle g_{\bm{\alpha}}\mid\bm{\alpha}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{\alpha}\leq\bm{\beta}_{0}\rangle$.
Since $\psi$ has the inverse element $\psi(\bm{t})^{-1}\in R\llbracket\bm{t}\rrbracket$, we have
$\psi(\bm{t})^{-1}\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r}}f_{\bm{n}}\bm{t}^{\bm{n}}=\sum_{\bm{\alpha}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{\alpha}\leq\bm{\beta}_{0}}g_{\bm{\alpha}}\bm{t}^{\bm{\alpha}}.$
By expanding the left-hand side and comparing the coefficients of $\bm{t}^{\bm{\alpha}}$ with $\bm{\alpha}\leq\bm{\beta}_{0}$ on both sides of the equation, we conclude that $g_{\bm{\alpha}}\in\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{n}\leq\bm{\beta}_{0}\rangle$ for all $\bm{\alpha}\leq\bm{\beta}_{0}$. Thus, we have $\langle g_{\bm{\alpha}}\mid\bm{\alpha}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{\alpha}\leq\bm{\beta}_{0}\rangle\subset\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{n}\leq\bm{\beta}_{0}\rangle$.
Summing up the above,
$\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r}\rangle\subset\langle g_{\bm{\alpha}}\mid\bm{\alpha}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{\alpha}\leq\bm{\beta}_{0}\rangle\subset\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{r},\bm{n}\leq\bm{\beta}_{0}\rangle.$
This proves the assertion. ∎
Example 2.2.
Let $f_{n}=xy^{2n+1}+zw^{2n+1}\in{\mathbb{Q}}[x,y,z,w]$ for $n\in{\mathbb{Z}}_{\geq 0}$. Then
$\sum_{n=0}^{\infty}f_{n}t^{n}$ $=xy\sum_{n=0}^{\infty}(y^{2}t)^{n}+zw\sum_{n=0}^{\infty}(w^{2}t)^{n}=\frac{xy}{1-y^{2}t}+\frac{zw}{1-w^{2}t}$
$=\frac{(xy+zw)-(xw+yz)ywt}{(1-y^{2}t)(1-w^{2}t)}.$
Since $\frac{1}{(1-y^{2}t)(1-w^{2}t)}$ is invertible, $\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle$ is generated by $f_{0},f_{1}$ by Lemma 2.1. This ideal appears as the defining ideal for the set of true parameters in a specific probabilistic model [Watanabe].
Definition 2.3.
Let ${\mathbb{Q}}[w]$ be a polynomial ring, and $k\in{\mathbb{Z}}_{\geq 0}$.
$\binom{w}{k}=\frac{w(w-1)\dots(w-k+1)}{k!}\in{\mathbb{Q}}[w]$
is called the binomial coefficient polynomial.
For $n,k\in{\mathbb{Z}}_{\geq 0}$, the binomial coefficient $\binom{n}{k}$ represents the number of ways to choose $k$ elements from a set of $n$ elements without regarding the order of selection, and thus $\binom{n}{k}\in{\mathbb{Z}}_{\geq 0}$. It is well-known that the generating functions of binomial coefficients are rational functions.
Theorem 2.4.
Let $k\in{\mathbb{Z}}_{\geq 0}$. For any commutative ring $R$,
$\sum_{n=0}^{\infty}\binom{n}{k}t^{n}=\dfrac{t^{k}}{(1-t)^{k+1}}$
holds in $R\llbracket t\rrbracket$.
Proof.
It suffices to prove the case where $R={\mathbb{Z}}$.
$\sum_{n=0}^{\infty}\binom{n}{k}t^{n}$ $=\dfrac{t^{k}}{k!}\sum_{n=0}^{\infty}n\cdot(n-1)\cdot(n-k+1)t^{n-k}$
$=\dfrac{t^{k}}{k!}\dfrac{d^{k}}{dt^{k}}\sum_{n=0}^{\infty}t^{n}=\dfrac{t^{k}}{k!}\dfrac{d^{k}}{dt^{k}}\dfrac{1}{1-t}=\dfrac{t^{k}}{(1-t)^{k+1}}.$
We are now ready to prove the main theorem of this paper.
Theorem 2.5.
Fix positive integers $p,r$, and $N$. Let $R$ be a commutative ring, and $\bm{x}=(x_{1},\dots,x_{r})\in R^{r}$. Let $c_{k}(\bm{n})\in R[\bm{n}]=R[n_{1},\dots,n_{p}]$ be polynomials in the parameters $\bm{n}=(n_{1},\dots,n_{p})$, and $\bm{A}^{(k)}=(\alpha^{(k)}_{ij})\in{\mathbb{Z}}_{\geq 0}^{r\times p},\bm{\beta}^{(k)}=(\beta^{(k)}_{i})\in{\mathbb{Z}}^{r}_{\geq 0}$ for $1\leq k\leq N$. Let
$f_{\bm{n}}=\sum_{k=0}^{N}c_{k}(\bm{n})\bm{x}^{\bm{A}^{(k)}\bm{n}+\bm{\beta}^{(k)}},~{}~{}\ell_{i}=\sum_{k=1}^{N}(\mathop{\mathrm{deg}}\nolimits_{n_{i}}c_{k}+1)-1,$
and $\bm{\ell}=(\ell_{1},\dots,\ell_{p})$.
Then the ideal $\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}^{r}_{\geq 0}\rangle\subset R$ is generated by $\{f_{\bm{n}}\mid\bm{n}\leq\bm{\ell}\}$.
Proof.
Let $\bm{\alpha}_{j}^{(k)}=(\alpha^{(k)}_{1j},\dots,\alpha^{(k)}_{rj})$. Then
$\bm{x}^{\bm{A}^{(k)}\bm{n}+\bm{\beta}^{(k)}}=\bm{x}^{\bm{\beta}^{(k)}}\prod_{j=1}^{p}\left(\bm{x}^{\bm{\alpha}_{j}^{(k)}}\right)^{n_{j}}.$
For each $k$, the coefficient polynomial $c_{k}(\bm{n})$ can be written as an $R$-linear summation of binomial coefficient polynomials:
$c_{k}(\bm{n})=\sum_{\begin{subarray}{c}\bm{m}=({m}_{1},\dots,{m}_{p})\in{\mathbb{Z}}_{\geq 0}^{p}\\ \forall j,0\leq{m}_{j}\leq\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}\end{subarray}}\gamma_{\bm{m}}\prod_{j=1}^{p}\binom{n_{j}}{{m}_{j}}$
for some $\gamma_{\bm{m}}\in R$. By Lemma 2.1
$\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}}\prod_{j=1}^{p}\binom{n_{j}}{{m}_{j}}\bm{x}^{\bm{A}^{(k)}\bm{n}+\bm{\beta}^{(k)}}\bm{t}^{\bm{n}}$ $=\bm{x}^{\bm{\beta}^{(k)}}\prod_{j=1}^{p}\left\{\sum_{n_{j}=0}^{\infty}\binom{n_{j}}{{m}_{j}}\left(\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{n_{j}}\right\}$
$=\frac{\bm{x}^{\bm{\beta}^{(k)}}\bm{t}^{\bm{m}}}{\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{{m}_{j}+1}}$
Hence
$\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}}f_{\bm{n}}\bm{t}^{\bm{n}}$ $=\sum_{k=1}^{N}\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}}c_{k}(\bm{n})\bm{x}^{\bm{A}^{(k)}\bm{n}+\bm{\beta}^{(k)}}\bm{t}^{\bm{n}}$
$=\sum_{k=1}^{N}\sum_{\begin{subarray}{c}\bm{m}=({m}_{1},\dots,{m}_{p})\in{\mathbb{Z}}_{\geq 0}^{p}\\ \forall j,0\leq{m}_{j}\leq\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}\end{subarray}}\frac{\gamma_{\bm{m}}\bm{x}^{\bm{\beta}^{(k)}}\bm{t}^{\bm{m}}}{\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{{m}_{j}+1}}.$
By putting all terms over a common denominator, we have
$\sum_{\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}}f_{\bm{n}}\bm{t}^{\bm{n}}=\frac{g(\bm{t})}{\prod_{k=1}^{N}\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}+1}},$
where the numerator $g(\bm{t})$ satisfies
$\mathop{\mathrm{deg}}\nolimits_{t_{j}}g(\bm{t})<\mathop{\mathrm{deg}}\nolimits_{t_{j}}\prod_{k=1}^{N}\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}+1}=\sum_{j=1}^{p}\left(\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}+1\right).$
This shows that any term $\bm{t}^{\bm{m}}$ appearing in $g(\bm{t})$ with non-zero coefficient satisfies $\bm{m}\leq\bm{\ell}$. Since the denominator $\prod_{k=1}^{N}\prod_{j=1}^{p}\left(1-\bm{x}^{\bm{\alpha}_{j}^{(k)}}t_{j}\right)^{\mathop{\mathrm{deg}}\nolimits_{n_{j}}c_{k}+1}$ is invertible in $R\llbracket\bm{t}\rrbracket$, the ideal $\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}^{r}_{\geq 0}\rangle\subset R$ is generated by $\{f_{\bm{n}}\mid\bm{n}\leq\bm{\ell}\}$ by Lemma 2.1. ∎
Example 2.6.
Let $R={\mathbb{Q}}[x,y,z]$ and
$f_{n}$ $=x^{2n+1}+y^{2n+1}+z^{n+2}+2^{n+2}x^{2}$
for $n\in{\mathbb{Z}}_{\geq 0}$. Since degree of the coefficients $1,1,1,1$ are all 00, we have $\ell=4-1=3$. By Theorem 2.5, $\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle$ is generated by $f_{0},\dots,f_{3}$.
Example 2.7.
Let $R={\mathbb{Q}}[x_{1},x_{2},x_{3},y_{1},y_{2},y_{3}]$ and
$f_{n}=(n+1)x_{1}y_{1}^{2n+1}+x_{2}y_{2}^{n+3}+(n^{2}+1)x_{3}y_{3}^{n+2}$
for $n\in{\mathbb{Z}}_{\geq 0}$. Since the degree of the coefficients $n+1,1,n^{2}+1$ are $1,0,2$ respectively, we have $\ell=(1+1)+(0+1)+(2+1)-1=5$. Thus $\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle$ is generated by $f_{0},\dots,f_{5}$ by Theorem 2.5.
Example 2.8.
Let $R={\mathbb{Q}}[x,y,z]$ and
$f_{n,m}=(n^{2}-m)x^{n+m+2}+(nm^{4}+m^{2})x^{2n+m+1}z^{m+n+3}+(n^{5}+n^{2})y^{4n+m+2}\in R.$
Then the degree of the coefficients $n^{2}-m,nm^{4}+m^{2},n^{5}+n^{2}$ in $n$ (resp. $m$) are $2,1,5$ (resp. $1,4,0$). Thus
$\ell_{1}$ $=(2+1)+(1+1)+(5+1)-1=10,$
$\ell_{2}$ $=(1+1)+(4+1)+(0+1)-1=7,$
and $\langle f_{n,m}\mid n,m\in{\mathbb{Z}}_{\geq 0}\rangle$ is generated by $\{f_{n,m}\mid 0\leq n\leq 10,0\leq m\leq 7\}$ by Theorem 2.5.
We remark that a generating set obtained by Theorem 2.5 is not always minimal.
Example 2.9.
Let $f_{n}=x^{2n+1}+(n^{100}+1)y^{n+1}$ for $n\in{\mathbb{Z}}_{\geq 0}$. The generating set of $\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle$ obtained by Theorem 2.5 is $f_{0},\dots,f_{101}$. On the other hand, $\langle f_{0},f_{1},f_{2}\rangle=\langle x+y,y^{2}\rangle$, and furthermore, it can be easily shown that $f_{n}\in\langle x+y,y^{2}\rangle$ for all $n\in{\mathbb{Z}}_{\geq 0}$.
3. The case of higher degree exponents
In this section, we consider the case where the exponents are polynomials in parameters of degree $\geq 2$. Since $\sum_{n=0}^{\infty}t^{n^{d}}$ is not a rational function for $d\geq 2$, we cannot apply the method used in the proof of Theorem 2.5 in this cases.
We will propose an approach to higher degree exponents case using Gröbner basis ([Buchberger][CLO]) and Theorem 2.5. By using Gröbner bases and the division algorithm, the ideal membership problem for finitely generated ideals can be solved (see [CLO] Section 2.8). The idea is as follows: Let $\bm{n}=(n_{1},\dots,n_{p})$ and $\bm{m}=(m_{1},\dots,m_{q})$ be parameter vectors, and let $\varphi_{j}(\bm{m})\in{\mathbb{Z}}[\bm{m}]$ for $1\leq j\leq q$. Let
$g_{\bm{m}}=\sum_{k=0}^{N}c_{k}(m_{1},\dots,m_{q})\prod_{i=1}^{r}x_{i}^{\sum_{j=1}^{q}\alpha^{(k)}_{ij}m_{j}+\beta^{(k)}_{i}}\in R[x_{1},\dots,x_{r}],$
$f_{\bm{n}}=g_{(\varphi_{1}(\bm{n}),\dots,\varphi_{q}(\bm{n}))}=\sum_{k=0}^{N}c_{k}(\varphi_{1}(\bm{n}),\dots,\varphi_{q}(\bm{n}))\prod_{i=1}^{r}x_{i}^{\sum_{j=1}^{q}\alpha^{(k)}_{ij}\varphi_{j}(\bm{n})+\beta^{(k)}_{i}},$
and $I=\langle f_{\bm{n}}\mid\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}\rangle$. We consider a finite generating set of $I$. Fix a sutable $\bm{k}_{0}\in{\mathbb{Z}}_{\geq 0}^{r}$ and let $J=\langle f_{\bm{n}}\mid\bm{n}\leq\bm{k}_{0}\rangle$. Then take a sufficient large $\bm{k}_{1}\in{\mathbb{Z}}_{\geq 0}^{r}$ with $\bm{k}_{0}<\bm{k}_{1}$, and compute the list
$L=\{\bm{m}\in{\mathbb{Z}}_{\geq 0}^{q}\mid\bm{m}\leq\bm{k}_{1},~{}g_{\bm{m}}\in J\}$
by using Gröbner bases. It sometimes happens that the list $L$ contain $\bm{m}$’s which are not written as $\bm{m}=(\varphi_{1}(\bm{n}),\dots,\varphi_{q}(\bm{n}))$ for any $\bm{n}\in{\mathbb{Z}}_{\geq 0}^{p}$. If we are fortunate, we can conclude that $f_{\bm{n}}\in J$ for all $\bm{n}$ using Theorem 2.5, which proves $I=J$.
Example 3.1.
Let $R={\mathbb{Q}}[x,y,z,w]$, and
$g_{n}$ $=xy^{n+1}+z^{n+1}+w^{n+1},$
$f_{n}=g_{n^{3}}$ $=xy^{n^{3}+1}+z^{n^{3}+1}+w^{n^{3}+1}.$
Let $I_{1}=\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle\subset R$ and $J_{1}=\langle f_{0},f_{1},f_{2},f_{3}\rangle=\langle g_{0},g_{1},g_{8},g_{27}\rangle$. Using Gröbner basis, we can determine that $g_{28},g_{29}\in J_{1}$ even though $28,29$ are not cubic numbers.
On the other hand, by Theorem 2.5, $\langle g_{n+27}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle$ is generated by $g_{27},g_{28},g_{29}$. Hence $g_{n}\in J_{1}$ for all $n\geq 27$. In particular, $f_{n}=g_{n^{3}}\in J_{1}$ for all $n\geq 3$. In conclusion, we have $I_{1}=J_{1}$.
Example 3.2.
Let $R={\mathbb{Q}}[x,y,z,w]$, and
$g_{n}$ $=(2n^{2}+3)xy^{n+2}+(n+1)z^{n+2}+w^{n+2},$
$f_{n}=g_{n^{3}}$ $=(2n^{6}+3)xy^{n^{3}+2}+(n^{3}+1)z^{n^{3}+2}+w^{n^{3}+2}.$
Let $I_{2}=\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle\subset R$ and $J_{2}=\langle f_{0},f_{1},f_{2},f_{3}\rangle=\langle g_{0},g_{1},g_{8},g_{27}\rangle$. By using Gröbner basis, we can determine that $g_{n}\in J_{2}$ for $n=27,28,29,30,31,32$. By Theorem 2.5, $\langle g_{n+27}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle$ is generated by $g_{n+27}$ for $0\leq n\leq 5$. Hence $g_{n}\in J_{2}$ for all $n\geq 27$. In particular, $f_{n}=g_{n^{3}}\in J_{2}$ for all $n\geq 3$. In conclusion, we have $I_{2}=J_{2}$.
Example 3.3.
Let $R={\mathbb{Q}}[x,y,z,w]$, and
$f_{n}$ $=xy^{n+1}+z^{n+1}+w^{n+1},$
$g_{n}=f_{n^{2}}$ $=xy^{n^{2}+1}+z^{n^{2}+1}+w^{n^{2}+1}.$
$I_{3}=\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle\subset R$ and $J_{3}=\langle f_{0},f_{1},f_{2},f_{3}\rangle=\langle g_{0},g_{1},g_{2},g_{9}\rangle$. Figure 1 shows the values of $n$ that satisfy $g_{n}\in J_{3}$, computed using Gröbenr basis. By Theorem 2.5, $g_{9},g_{12},g_{15}\in J_{3}$ and $g_{10},g_{13},g_{16}\in J_{3}$ implies $g_{3n+9},g_{3n+10}\in J_{3}$ for all $n\in{\mathbb{Z}}_{\geq 0}$. Hence $g_{n}\in J_{3}$ for all $n\geq 9$ with $n\not\equiv 2\bmod 3$. Since $n^{2}\not\equiv 2\bmod 3$ for all $n$, we conclude that $f_{n}\in J_{3}$ for all $n\geq 3$. In conclusion, we have $I_{3}=J_{3}$.
Figure 1. Plot of $n$ satisfying $g_{n}\in J_{3}$.
Example 3.4.
Let $R={\mathbb{Q}}[x,y,z,u,v]$, and
$g_{n_{1},n_{2}}$ $=x^{2n_{1}+3}+y^{2n_{2}+3}+u^{n_{1}+2}v^{n_{2}+2}$
$f_{n,m}=f_{n^{2}+m^{2},m^{3}}$ $=x^{2n^{2}+2m^{2}+3}+y^{2m^{3}+3}+u^{n^{2}+m^{2}+2}v^{m^{3}+2}.$
We will find a finite generating set of the ideal $I_{4}=\langle f_{n,m}\mid n,m\in{\mathbb{Z}}_{\geq 0}\rangle\subset R$. Let $J_{4}=\langle f_{n,m}\mid 0\leq n\leq 3,0\leq m\leq 2\rangle$. Figure 2 shows the pairs $(n_{1},n_{2})$ that satisfy $g_{n_{1},n_{2}}\in J_{4}$. Since $g_{n_{1}+1,n_{2}}\in J_{4}$ for $0\leq n_{1},n_{2}\leq 2$, we have $g_{n_{1}+1,n_{2}}\in J_{4}$ for all $n_{1},n_{2}\in{\mathbb{Z}}_{\geq 0}$ by Theorem 2.5. Since $n^{2}+m^{2}\geq 1$ for $(n,m)\neq(0,0)$, it follows that $f_{n,m}\in J_{4}$ for all $(n,m)\in{\mathbb{Z}}_{\geq 0}$. In conclusion, we have $I_{4}=J_{4}$.
Figure 2. Plot of $(n_{1},n_{2})$ satisfying $g_{n_{1},n_{2}}\in J_{4}$.
Example 3.5.
Let $R={\mathbb{Q}}[x,y,z,u,v]$, and
$g_{n_{1},n_{2}}$ $=x^{2n_{1}+3}+y^{2n_{2}+3}+u^{n_{1}+2}v^{n_{2}+2}$
$f_{n,m}=g_{n^{2}+m^{2},n+m^{3}}$ $=x^{2n^{2}+2m^{2}+3}+y^{2m^{3}+n+3}+u^{n^{2}+m^{2}+2}v^{m^{3}+n+2}.$
We will find a finite generating set of the ideal $I_{5}=\langle g_{n,m}\mid n,m\in{\mathbb{Z}}_{\geq 0}\rangle\subset R$. Let
$J_{5}$ $=\langle f_{n,m}\mid 0\leq n\leq 3,0\leq m\leq 2\rangle$
$=\langle g_{0,0},g_{1,1},g_{2,2},g_{4,2},g_{4,8},g_{5,3},g_{5,9},g_{8,10},g_{9,3},g_{10,4},g_{13,11}\rangle.$
Figure 2 shows the pairs $(n_{1},n_{2})$ that satisfy $g_{n_{1},n_{2}}\in J_{5}$.
By Theorem 2.5, it follows that $g_{2n_{1}+3,2n_{2}+3},g_{2n_{1}+4,2n_{2}+4}\in J_{5}$ for all $n_{1},n_{2}\in{\mathbb{Z}}_{\geq 0}$. Thus $g_{n_{1},n_{2}}\in J_{5}$ for all $n_{1},n_{2}\in{\mathbb{Z}}_{\geq 0}$ satisfying $n_{1}\equiv n_{2}\bmod 2$ and $(n_{1},n_{2})\geq(3,3)$.
Since $n^{2}+m^{2}\equiv m^{3}+n\bmod 2$ and $(n^{2}+m^{2},m^{3}+n)\geq(3,3)$ for all $(n,m)\in{\mathbb{Z}}_{\geq 0}^{2}$ with the exceptions of $(n,m)=(0,0),(0,1),(1,0),(1,1),(2,0)$, it follows that $f_{n,m}\in J_{5}$ for all $n,m\in\mathbb{Z}_{\geq 0}$. In conclusion, we have $I_{5}=J_{5}$.
Figure 3. Plot of $(n_{1},n_{2})$ satisfying $g_{n_{1},n_{2}}\in J_{5}$.
Example 3.6.
$g_{n_{1},n_{2}}$ $=x^{n_{1}+1}y+z^{n_{2}+1}+w^{n_{1}+1},$
$f_{n}=g_{n^{2},n}$ $=x^{n^{2}+1}y+z^{n+1}+w^{n^{2}+1}.$
Let $I_{6}=\langle f_{n}\mid n\in{\mathbb{Z}}_{\geq 0}\rangle\subset R$ and $J_{6}=\langle f_{0},f_{1},f_{2},f_{3},f_{4}\rangle$. Figure 4 shows the pairs $(n_{1},n_{2})$ that satisfy $g_{n_{1},n_{2}}\in J_{6}$ and the curve $(t^{2},t),~{}t\geq 0$. Since $g_{6m_{1}+7,2m_{2}+3}\in J_{6}$ for $(m_{1},m_{2})\leq(2,2)$, we have $g_{6m_{1}+7,2m_{2}+3}\in J_{6}$ for all $(m_{1},m_{2})\in{\mathbb{Z}}_{\geq 0}^{2}$ by Theorem 2.5. Similarly, we have
$g_{6m_{1}+9,2m_{2}+3},~{}g_{6m_{1}+10,2m_{2}+4},~{}g_{6m_{1}+12,2m_{2}+4}\in J_{6}$
for all $(m_{1},m_{2})\in{\mathbb{Z}}_{\geq 0}^{2}$. This indicates that $g_{n_{1},n_{2}}\in J_{6}$ for all $(n_{1},n_{2})\in{\mathbb{Z}}_{\geq 0}^{2}$ satisfying the following conditions
$(n_{1},n_{2})\geq(7,3),~{}~{}n_{1}\equiv n_{2}\bmod 2,~{}~{}n_{1}\not\equiv 2\bmod 3.$
Since $n^{2}\not\equiv 2\bmod 3$, we have $f_{n}\in J_{6}$ for all $n\geq 3$. In conclusion, we have $I_{6}=J_{6}$.
Figure 4. Plot of $(n_{1},n_{2})$ satisfying $g_{n_{1},n_{2}}\in J_{6}$, and the curve $(t^{2},t),t\geq 0$.